4- How many gram of Magnesium is in 5.6 gram of Mg(OH)2?
Answer #1

Molar mas = 2 Qin Molar mass of mg (OH) = 249 + (16+1)x2 g Our 8m = 24g + 17x2 g = Sog Thus so g of Mg (OH), contain = 24 g m